A big olive (m=0.11 kg) lies of an xy coordinate
Answer #1

Given that,

mo = 0.11 kg ; mn = 0.84 kg ; (1.2,4.0)m

Fo = 3.2 i + 4.9 j and Fn = -2.7 i - 3.6j

lFol = 5.6 N ; lFnl = 4.5 N

We know that, F = m a => a = F/m

ao = Fo/mo = 5.6/0.11 = 51 m/s2

an = Fn/mn = 4.5/0.84 = 5.36 m/s2

The intial velocity of COM is 0. From equation of motion we know that,

S = ut + 1/2 a t2

Do = 1/2 x ao t2

Do = 0.5 x 51 x 4.12 = 428.66 m

Dn = 1/2 an t2

Dn = 0.5 x 5.36 x 4.12 = 45.05 m